Grade 12Physical Chemistry

wen ethyl alcolhol and acetic acid mixed together in equimolar proportions,equilibrium is attained when two-third of the acid and alcohol are consumd.the euilibrium const of reaction is.?

Shriya Mehrotra
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
First write the reaction:
C2H5OH......+......CH3COOH......................=.CH3COOC2H5.....+.....H2O
a mole...................a mole.......................................0........................0......(initially)
a-x......................a-x............................................x.................x..(At equilibrium)
Here equimolar of both the reactants are mixed. Let the number of mole of reactant is a, V is the volume of the solution and x is mole reacted till equilibrium is reached.
according to the question, x=2/3 of a=(2/3)*a
Hence, a-x=a-(2/3)a=a/3
[Here I have written the amount of water formed and it will also apear in the expression of equilibrium.]
So
K=\frac{[H2O]*[CH3COOC2H5]}{[CH3COOH]*[C2H5OH]}
K=\frac{x/V*x/V}{(a-x)/V*(a-x)/V}
Here put value of x=2a/3 and calculate the value of K